# From Long Videos to Engaging Clips: A Human-Inspired Video Editing Framework with Multimodal Narrative Understanding

### v1

@@ 4.2 Statistics and Comparison
4.3 Evaluation Metrics
Early evaluation metrics in this field typically focused on computing precision and recall between the shots or clips of the edited video and a reference video Pardo et al. (2021); Hu et al. (2023). Considering that static metrics are often insufficient to capture the nuanced qualities of creative editing, we introduce a suite of quantitative metrics designed to evaluate both the edited videos and the editing methods from multiple perspectives. While some of these metrics require human judgment for evaluation, their definitions are clear and well-structured, which helps promote objectivity and reproducibility in assessment.
Diversity. Diversity is used to evaluate the creativity and variability of the generated outputs. Given the same raw footage, suppose the editing method produces $n$ edited videos $(V_{1},V_{2},...,V_{n})$. The diversity is measured based on pairwise dissimilarity, defined using the average inverse Intersection-over-Union (IoU) score:
$=1-\frac{2}{n(n-1)}\sum_{i<j}^{n}\frac{\left|V_{i}\cap V_{j}\right|}{\left|V_{i}\cup V_{j}\right|}$ (2)
Smoothness. We observe that automatic editing systems often suffer from abrupt or unnatural transitions. We define a metric to assess the smoothness of editing video as the average uninterrupted viewing duration before encountering a jarring transition. Specifically:
where $\mathrm{Interruption}(V)$ denotes the number of perceptible disruptions in the video.
$\mathrm{VEI}=\mathrm{Engagement}\times\mathrm{Smoothness}$ (5)
In addition to general-purpose metrics, we propose two additional metrics specifically for advertising scenarios.
Hook Rate. This metric estimates the likelihood that viewers will pause to watch the video upon encountering it, reflecting the attractiveness of the video’s opening.
Suspense Rate. Suspense Rate measures the degree to which the ending of the video instills curiosity or tension that motivates continued viewing. It reflects the model’s ability to craft compelling endings that drive viewer retention.
5 Experiments
5.1 Experiment Settings
Data. We sample 220 episodes from the DramaAD dataset and select several complete storylines from these videos, which typically span approximately 10 episodes and last 20 to 30 minutes. These videos will serve as the raw footage for the editing process.
@@ 5.1 Experiment Settings
    End2end editing method based on ASR-transcribed dialogue, which is a commonly used editing method. For fairness, we replace the prompts with those suitable for short dramas.
@@
Additionally, human-edited videos in the dataset will serve as the golden standard for editing reference.
Ablation Study Settings. We conduct ablation studies to analyze the contribution of each component in our framework.
a. w/o highlight detection We skip the highlight detection and directly select the opening and ending from all scenes.
b. w/o opening&ending selection We regard all starting and ending scenes as selected candidates.
c. w/o content pruning All clips from the opening to the ending scenes are kept without filtering.
5.2 Main Results
As shown in Table 2, our model outperforms all end-to-end baselines in terms of smoothness, achieving a higher overall viewing quality index while maintaining comparable levels of engagement. This improvement is largely attributed to our use of scene-level segmentation, which avoids interrupting character dialogues and actions. A similar trend is observed when replacing dialogue transcripts with scene narratives: even end-to-end methods benefit from improved smoothness under this setting.
When using scene narratives as input, our method also demonstrates greater diversity compared to end-to-end approaches. In contrast, end-to-end methods based on dialogue transcripts tend to produce more fragmented clips, which increases diversity but significantly compromises smoothness and overall viewing experience.
Furthermore, compared to directly predicting the final editing script, our step-by-step decomposition strategy leads to better performance in advertisement scenarios.
5.3 Ablation Analysis
Ablation studies reveal that removing highlight detection increases diversity but leads to a decrease in engagement. The increase in diversity stems from the model’s ability to freely select editing ranges across all scenes. However, without explicit highlight detection, the resulting clips are less likely to contain sufficiently engaging moments, ultimately compromising viewer engagement.
If the opening and ending scenes are not filtered, the edited video may begin or end at arbitrary points, making it less effective at capturing the viewer’s attention and sustaining engagement. This can significantly reduce user retention and lead to a sharp decline in key advertising metrics.
Omitting scene pruning leads to a lower proportion of engaging content in the final video, but it helps reduce abrupt visual transitions, thereby improving smoothness.  [^3] However, since no additional processing is applied to the ending effects when stitching multiple episodes, abrupt transitions may still occur occasionally.
@@ 6 Conclusion
In this work, we present a novel human-inspired framework for automatic video editing, addressing key limitations of existing editing approaches. By integrating multimodal narrative understanding and decomposing the editing process into intuitive sub-tasks, our method achieves superior editing quality. To further advance research in this area, we introduce DramaAD, the first short drama dataset with reference advertisement edits, providing valuable resources for both academia and industry. Experimental results demonstrate that our approach significantly narrows the gap between automated and human editing.
@@ A.2 Comprehensive Caption
Table 4 shows the details of the scene description we obtained using MLLM as well as multidimensional information
@@ A.6 End2End Editing
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 5.1 Experiment Settings
Data. We sample 220 episodes from DramaAD dataset and select several complete storylines from these videos, which typically span approximately 10 episodes and last 20 to 30 minutes. These videos will serve as the raw footage for the editing process.
@@
    End2end editing method based on ASR-transcribed dialogue, which is a commonly used editing method. For fairness, we replace the prompts with the same requirements.
@@
  • (3)
    Additionally, human-edited videos in the dataset will serve as the golden editing reference.
Ablation Study Settings. We conduct ablation studies to analyze the contribution of the components in our framework.
a. w/o highlight localization We directly view all scenes as starting and ending candidates.
b. w/o boundary selection We randomly select starting and ending scenes from candidates.
c. w/o pruning All clips from the opening to ending scenes are kept without filtering.
5.2 Experimental Results
Main Results. As shown in Table 2, our proposed method outperforms end-to-end editing approaches in both general and advertisement scenarios.
ASR-based methods yield higher diversity, likely due to richer timestamp information in conversational data, but result in lower fluency, limiting their industrial applicability. Comparing the two end-to-end methods, our video understanding framework improves fluency through semantic-based shot aggregation, albeit at the expense of some diversity.
Ablation Analysis. Ablation studies reveal that removing highlight detection increases diversity but lowers the engagement ratio. The increase in diversity stems from the model’s ability to freely select editing ranges across all scenes. Randomly selecting opening and ending scenes significantly weakens advertisement effectiveness and slightly decreases engagement. Omitting scene pruning leads to a lower proportion of engaging content in the final video, but it reduces abrupt visual transitions, thereby improving smoothness.  [^4] However, since no additional processing is applied to the ending effects when stitching multiple episodes, abrupt transitions may still occur occasionally.
@@ 6 Conclusion
In this work, we present a novel human-inspired framework for automatic video editing, addressing key limitations of existing ASR-based approaches. By integrating multimodal narrative understanding and decomposing the editing process into intuitive sub-tasks, our method achieves superior editing quality with improved interpretability. To further advance research in this area, we introduce DramaAD, the first large-scale short drama dataset with reference advertisement edits, providing valuable resources for both academia and industry. Experimental results demonstrate that our approach significantly narrows the gap between automated and human editing. We hope our work will advance research in the field of automatic video editing.
@@
Acknowledgments
This work was supported in part by the grants from National Natural Science Foundation of China (No.62222213, U22B2059).
@@ A.2 Comprehensive Caption
Table 4 shows the details of the scene description we obtained using MLLM as well as multidimensional information.
@@ A.6 End2End Editing
- Audience Engagement: The opening scene should immediately grab attention and draw viewers in. - Clarity: Avoid starting with scenes that rely heavily on prior plot context, such as ones mid-event, which might confuse new viewers. - Introduction Scenes: Scenes that introduce characters, setting, or plot premise can be good starting points. For Selecting End Scenes: You should analyze the plot and consider: - Relevance: Avoid ending on scenes unrelated to the highlight, as they may transition to a new, less exciting storyline. - Neutral Endings: If a scene doesn’t strongly affect the viewing experience, either way, it can still be chosen as an ending. - Suspense: Prefer ending on scenes that leave a sense of suspense, encouraging viewers to click to find out what happens next. - Complete Story Arc: It is also acceptable to end where a plot arc is fully wrapped up, giving a satisfying sense of closure while showcasing the highlight. {Scene narration of episode 1 $\sim$ T with tags}
@@
2. Do not remove the first or last scene, even if labeled as <Regular Scene>. 3. If there are no deletable scenes, simply output an empty list inside the <result> tag. 4. Maintain the original episode and scene_id numbers as provided. 5. Be cautious when removing scenes. Unnecessary deletions may affect the coherence of the storyline, negatively impacting the viewing experience. {Scene narration of episode 1 $\sim$ T with tags}
@@
- Audience Engagement: The opening scene should immediately grab attention and draw viewers in. - Clarity: Avoid starting with scenes that rely heavily on prior plot context, such as ones mid-event, which might confuse new viewers. - Introduction Scenes: Scenes that introduce characters, setting, or plot premise can be good starting points. For Selecting End Scenes: You should analyze the plot and consider: - Relevance: Avoid ending on scenes unrelated to the highlight, as they may transition to a new, less exciting storyline. - Neutral Endings: If a scene doesn’t strongly affect the viewing experience, either way, it can still be chosen as an ending. - Suspense: Prefer ending on scenes that leave a sense of suspense, encouraging viewers to click to find out what happens next. - Complete Story Arc: It is also acceptable to end where a plot arc is fully wrapped up, giving a satisfying sense of closure while showcasing the highlight. ## Output When generating the output, you should reflect on these requirements and devise an appropriate editing strategy. After careful consideration, select the scenes you believe are suitable for the final cut and include them in a list. Remember to use <result> as the delimiter and present the edited script in the following format: <result> [ {"episode":1, "scene_id": 0, "thought": Your justification for choosing this scene. } {"episode":1, "scene_id": 2, "thought": Your justification for choosing this scene. } ] </result> Explanation: - episode: episode number. - scene_id: scene number within that episode. - thought: your reasoning and decision for each scene. {Scene narration of episode 1 $\sim$ T}
@@
- Audience Engagement: The opening shots should immediately grab attention and draw viewers in. - Clarity: Avoid clips that rely heavily on prior plot context, such as ones mid-event, which might confuse new viewers. - Introduction Shots: Shots that introduce characters, setting, or plot premise can be good starting points. For Selecting Ending: You should analyze the plot and consider: - Relevance: Avoid endings that are unrelated to the highlight, as they may transition to a new, less exciting storyline. - Neutral Endings: If a shot doesn’t strongly affect the viewing experience, either way, it can still be chosen as an ending. - Suspense: Prefer endings that leave a sense of suspense, encouraging viewers to click to find out what happens next. - Complete Story Arc: It is also acceptable to end where a plot arc is fully wrapped up, giving a satisfying sense of closure while showcasing the highlight. ## Output In the output phase, you can consider these requirements and devise an editing strategy first. Finally, save the segments you deem suitable for editing into a list, and output your decision as a JSON list wrapped in <result> tags, like this: <result> [ {"episode":1, "start_time": 0, "end_time": 14, "thought": Your justification for choosing this clip. } {"episode":1, "start_time": 18, "end_time": 125, "thought": Your justification for choosing this clip. } ] </result> Explanation: - episode: episode number. - start_time: starting time of selected clip. - end_time: ending time of selected clip. - thought: your justification for choosing this clip. {ASR information of episode 1 $\sim$ T}